# Efficient Motion Sickness Assessment: Recreation of On-Road Driving on a Compact Test Track

### v1

@@ V-C Objective comparison of motion sickness assessment
For further investigation of the differences of these three cases, this section compares the objective motion sickness scores (MS metrics, Section III-D) based on the vehicle accelerations for the three cases (on-road, generated, and test track data). More specifically, Table III illustrates the MS metric calculated per acceleration signal (, with x, y, z, , , and ) and the percentage of difference from the on-road data. The gray rows in Table III illustrated the signals matched in the MPC framework. To generate the track motion, the vertical, roll and pitch dynamics were not considered. Here we also report these signals to quantify their impact on motion sickness.
According to Table III, similar MS metrics are obtained for all three cases illustrating the algorithms ability to replicate road sickness exposure. The on-road data’s MS score based on the longitudinal accelerations (MSx) shows a difference greater than 35 compared to the generated and test track data, respectively. This illustrates how the differences identified in Figures 8(a) and 7(a) translated in motion sickness levels. Furthermore, the MS score based on the lateral accelerations shows smaller differences especially when comparing the generated and the on-road data (3). This difference increases a bit due to the change of the acceleration by the experimental vehicle’s motion planner, when the generated path was applied on the test-track. The MS score based on the yaw accelerations has increased in the generated data by 34-38 compared to on-road. However, the amplitude is much lower compared to the MSx and MSy. Hence, the impact of the increase in the yaw accelerations is limited on the total MS score (MStotal). Similarly, the impact of the vertical accelerations is expected to be very small due to the low amplitude despite the significant difference between on-road and test-track (72). The MS score based on the roll and pitch accelerations is also significantly higher in the on-road than the test-track but their low amplitude marginally affects the MStotal.
@@
| MSx | 0.178 | 36 | 0.174 | 37 | 0.279 |
| MSy | 0.517 | 3 | 0.470 | 11 | 0.534 |
| MSz | x | x | 0.011 | 72 | 0.039 |
| MSr | 0.052 | 34 | 0.047 | 38 | 0.034 |
| MSϕ | x | x | 0.001 | 50 | 0.002 |
| MSθ | x | x | 0.001 | 50 | 0.002 |
| MStotal | 0.547 | 9 | 0.502 | 17 | 0.604 |
@@ VI-B Subjective assessment by human participants
Having objectively explored the efficiency of the algorithm, we subjectively assessed if the change of participants’ MISC levels over time in the two conditions differs. This is tested using the participants’ subjective input regarding motion sickness using MISC scores. During both conditions, the participants were engaged in watching sports videos as NDRT. According to the NDRT_counts, their focus during the ride was similar in both conditions and relatively high. More specifically, the participants identified on average 15.81 and 14.70 NDRT_counts out of 20 in total during the on-road and test-track condition.
@@
In addition to the efficient replication of the average motion sickness occurrence, the replication of the individual motion sickness occurrence is also critical. Hence, we opt to explore this through Figure 11, where we see a good correspondence at the extremes and at low levels of MISC. More specifically, five participants reported no sickness both on road and on track and five other participants reached the termination criterion (MISC=6) both on road and on track. Also, four and three participants respectively reported level 1 and 2 MISC in both conditions. The figure presents the correlation of the individual MISCmax between two conditions, fitting the data to a first order polynomial (y=a*x). The optimal fit is achieved with a = 0.746, with an adjusted = 0.45 (p0.001). Considering the high and the significance of the result (p0.001), this implies that the total individual MS occurrence (MISCmax) was replicated by 25 less in the test-track compared to the on-road.
@@ VII Limitation and Future Work
Test-track space limitations: The reproducability of the longitudinal accelerations in the test-track would have been higher if more space was available. However, even if this difference was captured by objective metrics, there was no difference between the participant’s responses during the test-track and the on-road scenarios. To further enhance the replication of on-road exposure more advanced MPC approaches can be used, such as the frequency-splitting MPC [30].
@@
Individual motion sickness occurrence: Even if our method was proven to accurately replicate the average motion sickness occurrence across multiple participants, the individual occurrence was replicated by a 25 decrease in the test-track compared to the on-road. This reduction was also captured by the objective metrics, but not from the subjective assessment. This could be studied and possibly improved by replacing the the MSDV with individual motion sickness models [31, 32] in the control modules of our motion planner as recently developed in the literature.
Engagement in NDRT: The participants during the experiment were instructed to focus on their non-driving related task. The NDRT ensured participants’ exposure to an environment similar to that of an AV and eliminated any additional sickening exposure from sensory conflicts arising through visual cues from the outside environment. However, according to the literature, visual sensory inputs are critical for the accumulation of MS [33]. Hence, further work by the authors is in progress to prove the generalization of the method for efficiently assessing motion sickness while not engaged in NDRTs.
Equipment requirements: Even if this method allows the efficient replicability of any driving scenario, an automated vehicle or steering and pedal robots are needed for the vehicle to follow the generated path.
@@ VIII Conclusions
In this work, a motion planner using model predictive control has been developed to replicate on-road sickness stimuli on a small test track. The algorithm outputs a reference velocity and trajectory profile for the vehicle to follow within the test-track boundaries. The effectiveness of the algorithm is proven by comparing objective metrics of motion sickness stimuli in time in terms of motion sickness dose value and frequency domain between the on-road, generated, and test track data. Furthermore, a within-subjects experiment with 47 human participants was conducted to test if the motion sickness occurrence did not differ between the two conditions. The results illustrated that our motion planner is able to replicate an on-road driving scenario in a small test track by maintaining highly similar motion sickness stimuli with lower speeds. This breakthrough proves that data collection can become more efficient and within a safer and controllable environment, allowing to speed up research into fundamentally understanding motion sickness in automated vehicles, modelling and prediction of motion sickness, and developing its mitigation strategies. Further work is in progress to exploit the collected data to unravel other differences that might rise between test-track and on-road testing.

### v2

@@ V-C Objective comparison of motion sickness assessment
For further investigation of the differences of these three cases, this section compares the objective motion sickness scores (MSDV metrics, Section III) based on the vehicle accelerations for the three cases (on-road, generated, and test track data). More specifically, Table III illustrates the MSDV metric calculated per acceleration signal (, with x, y, z, , , and ) and the percentage of difference from the on-road data. The gray rows in Table III illustrated the signals matched in the MPC framework. To generate the track motion, the vertical, roll and pitch dynamics were not considered. Here we also report these signals to quantify their impact on motion sickness.
According to Table III, similar MSDV metrics are obtained for all three cases illustrating the algorithms ability to effectively replicate road sickness exposure. The MSDV values are also representative and corresponding to the literature [31, 32, 33]. The on-road data’s MSDV score based on the longitudinal accelerations (MSx) shows a difference greater than 33 compared to the generated and test track data, respectively. This illustrates how the differences identified in Figures 7(a) and 6(a) translate in motion sickness levels. Furthermore, the MSDV score based on the lateral accelerations shows smaller differences especially when comparing the generated and the on-road data (6). This difference increases a bit due to the change of the acceleration by the experimental vehicle’s motion planner, when the generated path was applied on the test-track. The MSDV score based on the yaw accelerations has increased in the generated data by 38-50 compared to on-road. However, the amplitude is much lower compared to the MSDVx and MSDVy. Hence, the impact of the increase in the yaw accelerations is limited on the total MS score (MSDVt). Similarly, the impact of the vertical accelerations is expected to be very small due to the low amplitude despite the significant difference between on-road and test-track (72). The MSDV scores based on the roll and pitch accelerations are also significantly higher in the on-road than the test-track but their low amplitude marginally affects the MSDVt.
@@
| MSDVx [m/s1.5] | 7.28 | 33 | 6.83 | 37 | 10.83 |
| MSDVy [m/s1.5] | 21.94 | 6 | 18.48 | 11 | 20.70 |
| MSDVz [m/s1.5] | x | x | 0.45 | 71 | 1.52 |
| MSDVr [rad/s1.5] | 2.01 | 50 | 1.83 | 38 | 1.33 |
| MSDVϕ [rad/s1.5] | x | x | 0.03 | 48 | 0.07 |
| MSDVθ [rad/s1.5] | x | x | 0.04 | 49 | 0.08 |
| MSDVt [m/s1.5] | 23.19 | 1 | 19.79 | 17 | 23.44 |
@@ VI-B Subjective assessment by human participants
Having objectively explored the efficiency of the algorithm, we subjectively assessed if the change of participants’ MISC levels over time in the two conditions differs. This is tested using the participants’ subjective input regarding motion sickness using MISC scores. During both conditions, the participants were engaged in watching sports videos as NDRT. According to the NDRT_counts, their focus during the ride was similar in both conditions and relatively high. More specifically, the participants identified on average 15.81 and 14.70 NDRT_counts out of 16 in total during the on-road and test-track condition.
@@
In addition to the effective replication of the average motion sickness occurrence, the replication of the individual motion sickness occurrence is also critical. Hence, we opt to explore this through Figure 10, where we see a good correspondence at the extremes and at low levels of MISC. More specifically, three participants reported no sickness both on road and on track and five other participants reached the termination criterion (MISC=6) both on road and on track. Also, four and two participants reported level 1 and 2 MISC in both conditions, respectively . The figure presents the correlation of the individual MISCmax between two conditions, fitting the data to a first order polynomial (y=a*x). The optimal fit is achieved with a = 1.03, with an adjusted = 0.45 (p0.001).
@@ VII Limitation and Future Work
Test-track space limitations: The reproducability of the longitudinal accelerations in the test-track would have been higher if more space was available. However, even if this difference was captured by objective metrics, there was no significant difference between the participant’s responses during the test-track and the on-road scenarios. To further enhance the replication of on-road exposure more advanced MPC approaches can be used, such as the frequency-splitting MPC [32].
@@
Motion sickness occurrence: Our method was proven effective in replicating on-road driving scenarios with low MISC levels, but it is expected to be able to also replicate on-road driving scenarios with higher MISC levels.
Engagement in NDRT: The participants were instructed to focus on the non-driving related task of video watching. This ensured participants’ exposure to an environment similar to using an AV in an eyes-off-road mode. In practice, AV users will occasionally observe the road, and non-driving tasks may vary in visual and motor complexity, which have a modest but significant effect on MS [33]. Hence, further work by the authors is in progress to prove the generalization of the method for efficiently assessing motion sickness while not engaged in NDRTs and other conditions.
Equipment requirements: Even if this method allows the effective replicability of any driving scenario, an automated vehicle or steering and pedal robots are needed for the vehicle to follow the generated path.
@@ VIII Conclusions
In this work, a motion planner using model predictive control has been developed to replicate on-road sickness stimuli on a small test track. The algorithm outputs a reference velocity and trajectory profile for the vehicle to follow within the test-track boundaries. The effectiveness of the algorithm is proven by comparing objective metrics of motion sickness stimuli in time in terms of motion sickness dose value and frequency domain between the on-road, generated, and test track data. Furthermore, a within-subjects experiment with 47 human participants was conducted to test if the motion sickness occurrence did not differ between the two conditions. The results illustrated that our motion planner is able to replicate an on-road driving scenario in a small test track by maintaining highly similar motion sickness stimuli with lower speeds. This breakthrough proves that data collection can become more efficient and within a safer and controllable environment, allowing to speed up research into fundamentally understanding motion sickness in automated vehicles, modelling and prediction of motion sickness, and developing its mitigation strategies. This method enables testing of early stage prototype or unreleased developed stages without approval of driving on pubic roads, securing realistic results with regards to on-road driving. Further work is in progress to exploit the collected data to unravel other differences that might rise between test-track and on-road testing.